# Fine-grained Stateful Knowledge Exploration: Effective and Efficient Graph Retrieval with Large Language Models

### v3

@@ 3.5 Ending exploration and answering questions
IO prompt Brown et al. (2020) w/ChatGPT 8.0 7.3 10.0 7.5
CoT Wei et al. (2022b) w/ChatGPT 3.4 2.2 10.0 6.7
@@
Ours w/GLM4-9B-Chat 84.1 55.0 77.1 67.9
Ours w/ChatGPT 89.1 71.1 83.1 76.4
Ours w/GPT-4 94.4 74.4 84.2 77.8
4 Experiments
4.1 Experimental Settings
Datasets and Evaluation Metrics. To evaluate the performance of our proposed paradim, We selected two open-source knowledge graphs and one self-constructed graph as external knowledge bases for experimentation. The open-source knowledge graphs are MOOC Q&A Yu et al. (2020) and Freebase Bollacker et al. (2008). The graph that we constructed ourselves is not yet publicly available and is referred to in this paper as the agricultural knowledge graph.
Freebase is a large-scale, semi-structured database supported by Google, designed to collect and connect information about millions of entities and their relationships worldwide. It covers a wide range of topics such as people, places, books, movies, music, etc., with each entity having a unique identifier. We conduct experiments on three QA sets with Freebase as external knowledge base: WebQSP Yih et al. (2016) and WebQuestions Berant et al. (2013).
[…]
Baselines. We compared our approach with three baseline methods: standard prompting (IO prompt) Brown et al. (2020), chain of thought prompting (CoT prompt) Wei et al. (2022b) and ToG Sun et al. (2024). The knowledge base and Q&A sets in MOOC Q&A are highly interconnected, with most questions requiring knowledge retrieval from the knowledge base for answers. IO prompt and CoT prompt are two methods without external knowledge, and they are used to illustrate the proportion of questions requiring knowledge. ToG is a previous state-of-the-art method in knowledge base question-answering, used to validate the effectiveness of our knowledge retrieval approach.
Experiment Details. To ensure the reliability and reproducibility of the experiments, we set the temperature parameter to 0 for all LLMs. The two hyperparameters of ToG are set to their default values. We modified all prompts in ToG to be suitable for MOOC Q&A and the agricultural knowledge grpah. During the experiment, we observed that ToG would invoke LLM hundreds of times to answer certain questions, exceeding ten minutes. To mitigate the impact of these extreme cases, we terminated the exploration for all methods when it exceeded 30 LLM calls.
Table 2: Results for WebQuestions and WebQSP. To better illustrate whether these methods answer questions based on the inherent knowledge of LLMs or the retrieved knowledge, we merged the responses of these methods with the IO answers of LLMs and conducted additional comparisons. These are respectively represented as ToG + IO and Ours + IO. For each dataset, we conducted experiments on 300 data samples.
@@ 4.1 Experimental Settings
COT Wei et al. (2022b) w/Llama3-8B-Instruct 54.3 35.8 62.1 33.7
COT Wei et al. (2022b) w/GLM4-9B-Chat 54.8 34.3 63.0 32.5
@@
Ours w/Llama3-8B-Instruct 70.3 44.6 70.8 40.4
Ours w/GLM4-9B-Chat 66.3 38.8 62.5 36.4
Ours + IO w/Llama3-8B-Instruct 81.2 50.5 75.0 43.3
Ours + IO w/GLM4-9B-Chat 77.9 50.2 73.9 43.2
@@ 4.2 Performance Comparison
On the True/False sets, our method demonstrates outstanding performance, surpassing the previous state-of-the-art method ToG. Even the performance of our method based on GLM4-9B surpasses that of ToG based on GPT-4. This demonstrates the significant advantages of the new paradigm we propose over existing paradigms. The method we proposed showed improvements when different models such as GLM4-9B, ChatGPT, and GPT4 were used. This indicates that our method is very friendly to LLMs with fewer parameters, enabling better results at a lower cost, while still offering considerable room for improvement with better models.
@@
On the Query set, considering the partial match metric, in the case of GPT4, ToG’s results are only slightly worse than our method. However, when based on ChatGPT and GLM4-9B, ToG performs significantly worse than our method. This once again highlights the usability of our approach with LLMs that have smaller parameter sizes. Examining the complete match metric, our method consistently outperforms ToG across all scenarios. This indicates that while ToG can identify partial answers to questions, it struggles to comprehensively retrieve all possible answers to a given question. In contrast, our method excels in providing a comprehensive set of answers to the posed questions.
@@
The results of WebQuestions and WebQSP are shown in Table 2. Due to economic considerations and OpenAI’s policies, we did not test ChatGPT and GPT-4 for WebQSP and WebQuestions. From the baseline, we can observe that LLMs can answer a considerable portion of the questions solely based on their own knowledge, indicating that these questions are relatively familiar to LLMs. To better illustrate whether these methods answer questions based on the inherent knowledge of LLMs or the retrieved knowledge, we merged the responses of these methods with the IO answers of LLMs and conducted additional comparisons. These are respectively represented as ToG + IO and Ours + IO. It can be seen that even when using a massive external knowledge base like Freebase, our method still maintains a comprehensive advantage over ToG. It is worth noting that ToG requires manual provision of a topic entity, which means it can directly find a starting point in the knowledge base. Our method does not use manually provided topic entities. Comparing the result improvements of both methods after adding IO, it can be seen that IO provides a greater boost to the results of our method. This means that the results of our method are more often based on correctly retrieving knowledge from the knowledge base and answering questions, rather than relying on the inherent knowledge of the LLM.
The results of the agricultural graph are shown in Table 3. It can be seen that the method we proposed outperforms ToG on single-hop questions, but does not show an advantage on multi-hop questions. This agricultural dataset, which we constructed ourselves, has highly specialized questions, and multi-hop question-answering is relatively more complex. This may indicate that our method is not superior to existing methods when dealing with specialized and complex questions.
@@
In summary, the new method we proposed has obvious advantages over existing methods, with significantly higher accuracy and completeness of answers compared to existing methods. However, when encountering overly specialized and complex questions, our method performs on par with existing methods.
@@ 4.3 Computational Cost
In the knowledge base question-answering process, the majority of computational resources are dedicated to utilizing the Large Language Model, while the remaining computational consumption is relatively negligible. We calculate the average number of LLM calls during the question-answering process to illustrate the computational resource consumption. The results are shown in Table 4 and 5.
The results show that our method and ToG exhibit consistency in the number of LLM invocations across different question-answering sets. On more complex question sets, there are more LLM invocations. When encountering the same question-answering set, our method invokes the LLM significantly fewer times than ToG, roughly only one-third to half as many. We believe that this advantage is due to the stateful exploration we proposed, which can avoid redundant exploration.
Table 4: Average Number of Calls to LLMs On MOOC Q&A
@@
ToG Sun et al. (2024) w/ChatGLM4-9B 14.76 16.82 8.94
ToG Sun et al. (2024) w/ChatGPT 19.81 22.16 11.2
ToG Sun et al. (2024) w/GPT-4 22.6 26.17 8.06
Ours w/ChatGLM4-9B 6.64 9.32 4.71
Ours w/ChatGPT 5.86 8.18 5.58
Ours w/GPT-4 5.19 6.93 4.97
Table 5: Average Number of Calls to LLMs On the agricultural graph
Method 1-Hop Query Multi-Hop Query
ToG Sun et al. (2024) w/ChatGLM4-9B 11.64 16.2
Ours w/ChatGLM4-9B 5.72 9.45
Table 6: Ablation Studies on WebQuestions set. We conducted three experiments: removing the stateful exploration, removing the ending flag, and removing both.
Method partial match complete match avg LLM Calls
Ours 70.3 44.6 8.76
w/o Stateful exploration 65.7 44.3 9.06
w/o ending flag 59.4 42.6 9.12
w/o both 51.4 38.7 15.9
@@ 4.4 Ablation Studies
We conducted ablation studies and the results of ablation studies are shown in Table 6. We conducted three experiments: removing the stateful exploration, removing the termination flag, and removing both. Removing the ending flag also means that it is impossible to determine whether knowledge should be provided to the LLM or whether to directly ask the LLM a question. It can be seen that both stateful exploration and the ending flag contribute to reducing the number of LLM invocations, and when both parts are removed, the number of LLM invocations increases significantly. This is because both steps are based on fine-grained knowledge. Fine-grained knowledge, by mapping to the retrieved knowledge, naturally avoids redundant exploration.
When either of the two steps was removed, the question-answering results declined, with a more significant drop observed when the ending flag was removed. We analyzed that this is because the ending flag also serves as the basis for determining whether to provide knowledge to the LLM. After removing the ending flag, it becomes impossible to judge whether the retrieval is correct and complete, so the retrieved knowledge must be provided to the LLM every time. When retrieval fails, the erroneously retrieved knowledge interferes with the LLM’s response, leading to a decline in results.
@@ 5 Conclusion
In this paper, we propose a new paradigm of fine-grained stateful exploration. It aims to address the widespread issue of knowledge granularity mismatch in existing knowledge graph question-answering paradigms. We analyze that this problem leads to redundant exploration or omission of important information during the knowledge exploration process, ultimately resulting in increased computational consumption and degraded question-answering performance. In the new paradigm we propose, fine-grained information is mapped to the retrieved knowledge, effectively avoiding repetition and omissions. We conducted experiments based on three different knowledge graphs and corresponding multiple question-answering sets, and the results validated our design concepts. Our method significantly outperforms existing methods in terms of both question-answering results and computational efficiency. Additionally, the experiments demonstrate that our method still performs well with LLMs that have fewer parameters, showcasing great practical potential. Ablation studies also prove the effectiveness of each component of our design.

### v4

@@ 4.1 Experimental Settings
IO prompt Brown et al. (2020) w/GPT-3.5 8.0 7.3 10.0 7.5
CoT Wei et al. (2022) w/GPT-3.5 3.4 2.2 10.0 6.7
@@
PoG Chen et al. (2024) w/GLM4-9B-Chat 57.0 44.6 52.3 43.3
PoG Chen et al. (2024) w/GPT-3.5 67.5 52.9 73.9 54.1
PoG Chen et al. (2024) w/GPT-4 77.5 62.9 82.9 64.5
FiSKE(Ours) w/GLM4-9B-Chat 84.1 55.0 77.1 67.9
FiSKE(Ours) w/GPT-3.5 89.1 71.1 83.1 76.4
FiSKE(Ours) w/GPT-4 94.4 74.4 84.2 77.8
Experiment Details. To ensure the reliability and reproducibility of the experiments, we set the temperature parameter to 0 for all LLMs. The two hyperparameters of ToG are set to their default values. We modified all prompts for all baseline methods to be suitable for MOOC Q&A and the agricultural knowledge grpah. During the experiment, we observed that ToG would invoke LLM hundreds of times to answer certain questions, exceeding ten minutes. To mitigate the impact of these extreme cases, we terminated the exploration for all methods when it exceeded 30 LLM calls. We conduct experiments on one thousand question-answer pairs from each dataset.
Table 2: Results for WebQuestions and WebQSP. To better illustrate whether these methods answer questions based on the inherent knowledge of LLMs or the retrieved knowledge, we merged the responses of these methods with the IO answers of LLMs and conducted additional comparisons. These are respectively represented as ToG + IO, PoG + IO and FiSKE + IO.
@@
SC Wang et al. (2023a) w/Llama3-8B-Instruct 45.2 38.8 52.1 26.7
SC Wang et al. (2023a) w/GLM4-9B-Chat 54.4 34.9 59.6 32.1
COT Wei et al. (2022) w/Llama3-8B-Instruct 54.3 35.8 62.1 33.7
COT Wei et al. (2022) w/GLM4-9B-Chat 54.8 34.3 63.0 32.5
@@
PoG Chen et al. (2024) w/Llama3-8B-Instruct 62.3 39.4 60.2 35.2
PoG Chen et al. (2024) w/GLM4-9B-Chat 63.4 38.3 63.4 34.4
PoG w/GPT-3.5 - - 82.0 -
PoG + IO w/Llama3-8B-Instruct 70.9 42.1 70.8 40.3
PoG + IO w/GLM4-9B-Chat 73.6 45.2 74.4 41.6
StructGPT Jiang et al. (2023) w/GPT-3.5 - - 72.6 -
KB-BINDER Li et al. (2023a) w/GPT-3.5 - - 74.4 -
FiSKE w/Llama3-8B-Instruct 70.3 44.6 70.8 40.4
FiSKE w/GLM4-9B-Chat 66.3 38.8 62.5 36.4
FiSKE w/GPT-3.5 - - 83.1 49.1
FiSKE + IO w/Llama3-8B-Instruct 81.2 50.5 75.0 43.3
FiSKE + IO w/GLM4-9B-Chat 77.9 50.2 73.9 43.2
@@ 4.2 Performance Comparison
On the True/False sets, FiSKE demonstrates outstanding performance, surpassing the previous state-of-the-art methods. In 1-hop questions, even the performance of FiSKE based on GLM4-9B-Chat surpasses baseline methods based on GPT-4. In multi-hop questions, FiSKE still demonstrates significant advantages. This demonstrates the significant advantages of the new paradigm we propose over existing paradigms. FiSKE showed improvements when different models such as GLM4-9B-Chat, GPT-3.5, and GPT-4 were used. This indicates that FiSKE is very friendly to LLMs with fewer parameters, enabling better results at a lower cost, while still offering considerable room for improvement with better models.
@@
On the Query set, considering the partial match metric, in the case of GPT4, results of ToG and PoG are only slightly worse than FiSKE. However, when based on GPT-3.5 and GLM4-9B-Chat, ToG and PoG perform significantly worse than FiSKE. This once again highlights the usability of our approach with LLMs that have smaller parameter sizes. Examining the complete match metric, our method consistently outperforms existing methods across all scenarios. This indicates that while existing methods can identify partial answers to questions, it struggles to comprehensively retrieve all possible answers to a given question. In contrast, our method excels in providing a comprehensive set of answers to the posed questions.
@@
The results of WebQuestions and WebQSP are shown in Table 2. From the baseline, we can observe that LLMs can answer a considerable portion of the questions solely based on their own knowledge, indicating that these questions are relatively familiar to LLMs. To better illustrate whether these methods answer questions based on the inherent knowledge of LLMs or the retrieved knowledge, we merged the responses of these methods with the IO answers of LLMs and conducted additional comparisons. These are respectively represented as Method + IO.
It can be observed that when dealing with a large-scale external knowledge base like Freebase, our proposed FiSKE still maintains excellent performance, generally outperforming existing methods across the board. FiSKE consistently demonstrates superior performance under both partial and complete matching evaluation criteria, and FiSKE + IO demonstrates potential for substantial further improvement. This indicates that FiSKE’s significant advantage over existing methods stems primarily from its more accurate retrieval of knowledge from the knowledge graph.
The results of the agricultural graph are shown in Table 3. It can be seen that the method we proposed outperforms ToG on single-hop questions, but does not show an advantage on multi-hop questions. The agricultural knowledge graph we constructed exhibits high density, where the same set of entities can form different pathways through diverse relational connections. FiSKE’s suboptimal performance on multi-hop QA tasks likely stems from its current capability to only identify that all entities along these pathways are relevant to the clue, while lacking sufficient discriminative information to make precise selections among multiple candidate pathways. In our future work, we plan to extract richer contextual signals from the input questions to facilitate more precise path selection in the knowledge graph.
@@
In summary, FiSKE has obvious advantages over existing methods, with significantly higher accuracy and completeness of answers compared to existing methods. However, when encountering overly specialized and complex questions, FiSKE performs on par with existing methods.
@@ 4.3 Computational Cost
In the knowledge base question-answering process, the majority of computational resources are dedicated to utilizing the Large Language Model, while the remaining computational consumption is relatively negligible. We calculate the average number of LLM calls during the question-answering process to illustrate the computational resource consumption. The results are shown in Table 4.
We conducted experiments across different datasets and various base models. The results demonstrate that different question types significantly affect the number of LLM calls, while different base models also cause fluctuations in call frequency. However, overall, PoG requires significantly fewer LLM calls than ToG, and FiSKE requires even fewer calls than PoG. This clearly demonstrates FiSKE’s superior efficiency in knowledge retrieval, and its high performance can effectively facilitate the practical application of question-answering methods.
Table 4: Average Number of Calls to LLMs On MOOC Q&A.
@@
ToG Sun et al. (2024) w/GLM4-9B-Chat 14.76 16.82 8.94
ToG Sun et al. (2024) w/GPT-3.5 19.81 22.16 11.20
ToG Sun et al. (2024) w/GPT-4 22.60 26.17 8.06
PoG Chen et al. (2024) w/GLM4-9B-Chat 11.87 14.89 7.62
PoG Chen et al. (2024) w/GPT-3.5 14.53 15.71 7.29
PoG Chen et al. (2024) w/GPT-4 11.21 13.14 6.82
FiSKE w/GLM4-9B-Chat 6.64 9.32 4.71
FiSKE w/GPT-3.5 5.86 8.18 5.58
FiSKE w/GPT-4 5.19 6.93 4.97
4.4 Studies on Clue Extraction
We evaluate the impact of clue extraction on the overall method. Four different clue variants were established in our study, as shown in Table 5. (a) SP-Matched: Only clues that directly match a starting point in the knowledge graph are retained. (b) Clue-Trunc: We randomly select one clue (not starting point) to modify by deleting the original clue, then keeping only its first word (for multi-word clues) or repeating it (for single-word clues). Other clues remain unchanged. (c) Clue-Ext: Retains all original clues while adding their truncated variants to one randomly selected clue (same modification rules as (b)). (d) Noise-Add: Introduces irrelevant clues to measure resistance to noise.
Approximately 21% of QA pairs failed to locate starting points in the knowledge graph. For these cases, FiSKE defaults to CoT reasoning. When excluding these problematic pairs (as shown in SP-Matched results), accuracy improves at the cost of significantly higher computational overhead and processing time. The Clue-Trunc variant demonstrates reduced accuracy and computational costs because truncated clues often prematurely terminate knowledge exploration when failing to find valid paths. In Clue-Ext experiments, the additional clues increase exploration cycles, leading to substantially higher computational overhead while paradoxically decreasing accuracy. The irrelevant clues trigger unnecessary exploration cycles in the knowledge graph, all of which fail immediately due to semantic mismatch. While this still incurs computational overhead, the costs remain slightly lower than Clue-Ext because failed explorations terminate faster. All unsuccessful queries automatically default to CoT responses, resulting in accuracy identical to pure CoT performance.
Table 5: Experimental Results of Clue Extraction Variants on WebQSP (Llama3-8B-Instruct).
Dataset Method Partial Match Complete Match LLM Call Total Token LLM Time Total Time
WebQSP FiSKE 70.8 40.4 6.9 3814.2 49.5 57.7
SP-Matched 72.6 41.9 8.4 4819.1 61.6 72.0
Clue-Trunc 67.4 39.1 5.4 2583.0 31.1 39.0
Clue-Ext 64.1 36.2 8.4 4664.3 60.9 73.1
Noise-Add 62.1 33.7 7.9 4263.0 60.2 77.5
Table 6: Results for different backbone LLMs on WebQSP. Hybrid denotes a configuration where clue extraction is performed by GPT-3.5 while adaptive mapping strategy is executed on Llama3-8B-Instruct.
Method Partial Match Complete Match
FiSKE w/GPT-3.5 83.1 49.1
FiSKE w/Hybrid 80.6 46.5
FiSKE w/Llama3-8B-Instruct 70.8 40.4
4.5 Studies on Different Backbone LLMs
We conducted experiments on how FiSKE’s performance varies when implementing its Clue Extraction and Adaptive Mapping Strategy (AMS) modules with different LLMs, as shown in Table 6. When replacing the LLM used in the AMS phase from GPT-3.5 to Llama3-8B-Instruct, FiSKE’s performance showed a slight decline, indicating that AMS has relatively low requirements for LLM capability. However, when substituting the LLM used in the clue extraction phase from GPT-3.5 to Llama3-8B-Instruct, FiSKE’s performance decreased significantly. This primarily occurs because the correctness of clue extraction substantially impacts the method’s effectiveness, as further demonstrated in Section 4.4.
Table 7: Ablation Studies on WebQuestions set. SR represents the Stateful Record, AMS represents the Adaptive Mapping Strategy, and BAA represents Branch-Adaptive Answering.
Variant SR AMS BAA partial match complete match avg LLM Calls
FiSKE 70.3 44.6 8.76
w/o BAA × 58.4 35.6 8.76
w/o SR × 63.7 42.7 5.94
w/o SR+BAA × × 48.3 31.7 5.94
w/o AMS × 65.4 41.3 5.57
w/o AMS+BAA × × 42.6 24.1 5.57
w/o SR+AMS × × 61.2 40.0 6.24
w/o SR+AMS+BAA × × × 41.9 23.8 6.24
@@ 4.6 Ablation Studies
We conducted ablation studies and the results are shown in Table 7. We conducted ablation experiments on the three innovative components of FiSKE. In the Table 7, SR stands for Stateful Record, AMS represents the Adaptive Mapping Strategy, and BAA denotes Branch-Adaptive Answering. Specifically, when removing the Stateful Record, we still conducted experiments based on fine-grained clues but no longer recorded the mappings between clues and entities or relationships in the knowledge graph. When removing the Adaptive Mapping Strategy, we still performed relation mapping exploration and entity mapping exploration but enforced their mapping to the same clue rather than adaptively selecting one or multiple clues. When removing Branch-Adaptive Answering, we no longer determined whether to leverage the LLM’s own knowledge but uniformly passed the retrieved triples to the LLM. We not only tested the effects of removing each component individually but also experimented with the results of removing their combinations.
Overall, we can observe that each new component of FiSKE contributes to improving the overall performance of the method. Notably, Branch-Adaptive Answering has a particularly significant impact on the results. We argue that when the retrieved results are incomplete or even incorrect, directly feeding them to the LLM can substantially hinder its ability to correctly utilize its own knowledge for answering—especially for smaller-scale LLMs. Therefore, our strategy only provides the retrieved knowledge to the LLM after successfully completing all clue mappings, effectively mitigating this negative effect.
Another observation is that removing either Stateful Record or Adaptive Mapping Strategy reduces the number of LLM calls. Our analysis suggests that without these components, FiSKE tends to prematurely terminate exploration and instead directly query the LLM. While this decreases invocation frequency, it undermines effective knowledge retrieval from the knowledge graph. Consequently, when faced with questions beyond the LLM’s capability, performance degrades significantly.
Finally, there is an implicit yet noteworthy observation: Even when we simultaneously remove both Stateful Record and Adaptive Mapping Strategy—effectively reducing FiSKE’s knowledge exploration process to a level comparable to ToG—its performance still surpasses ToG. We attribute this advantage to our retention of fine-grained clues. Unlike ToG, which conducts knowledge exploration at the entire-question level, targeting fine-grained clues leads to more precise exploration results.
@@ 5 Conclusion
In this paper, we propose a new paradigm of fine-grained stateful exploration called FiSKE. It aims to address the widespread issue of knowledge granularity mismatch in existing knowledge graph question-answering paradigms. We analyze that this problem leads to redundant exploration or omission of important information during the knowledge exploration process, ultimately resulting in increased computational consumption and degraded question-answering performance. After decomposing questions into fine-grained clues, FiSKE maintains a stateful record to avoid redundant exploration or omissions, employs an adaptive mapping strategy to resolve ambiguity in clue-to-graph mappings, and automatically determines how explored knowledge paths are used for LLM augmentation. We conducted experiments based on three different knowledge graphs and corresponding multiple question-answering sets, and the results validated our design concepts. Our method significantly outperforms existing methods in terms of both question-answering results and computational efficiency. Additionally, the experiments demonstrate that our method still performs well with LLMs that have fewer parameters, showcasing great practical potential. Ablation studies also prove the effectiveness of each component of our design.
Appendix A Appendix
A.1 Prompt Templates
Our method employs four prompt templates to guide the LLM in performing four different tasks. Example prompts are shown in Table 8 and Fig. 3. FiSKE is not highly sensitive to the quality of the prompts. Our experiments include both Chinese and English datasets. We simply translated the prompts and modified the examples within them, without making any further specialized adjustments. For datasets in the same language, we used the same set of prompts. In Section 3.2, we use the fine-grained clue extraction prompt to obtain fine-grained clues. In Section 3.4, we apply the relation mapping prompt and the entity mapping prompt to complete the two steps of mapping exploration. Finally, in Section 3.5, we utilize the answering prompt to instruct the LLM to answer questions based on the provided knowledge.
Table 8: Examples of all LLM prompts used in this paper. These are designed for the Freebase knowledge base and can be appropriately adapted for other knowledge bases.
fine-grained clue extraction Please identify the entities in the sentence by taking into account the sentence’s syntactic structure. Below is an example. sentence: Which place is the madam satan located? entities: [’place’, ’madam satan’] Now answer with the format of the example above. sentence: {}
relation mapping Please retrieve relations that are related to any of the target information from the sentence and rate it on a scale of 0 to 10. Be brief and concise. sentence: Name the president of the country whose main spoken language was Brahui in 1980? target information: [’president’, ’country’, ’main spoken language’, ’Brahui’, ’1980’] relations: [’language.human_language.main_country’, ’language.human_language.human_language’, ’language.human_language.language_family’, ’language.human_language.iso_639_3_code’, ’base.rosetta.languoid.parent’, ’language.human_language.writing_system’, ’base.rosetta.languoid.languoid_class’] assess: - ’language.human_language.main_country’ is related to the country, so it’s a match, (9) score. - ’language.human_language.human_language’ is kind of related to the main spoken language, so it’s a match, (6) score. - other relations are not related to any information in the sentence, so they are not match, (0) score. sentence: {} target information: {} relations: {} assess:
entity mapping Please assess the relevance of each candidate entity to the information in the sentence and assign a score ranging from 0 to 10. The format of the candidate entity is (relation, entity). sentence: Which place is the madam satan located? information in the sentence: [’place’, ’madam satan’] candidate entity: [(’film.film.country’, ’the USA’), (’film.film.language’, ’English’)] assess: - (’film.film.country’, ’the USA’) is related to ’place’, so it’s a match, (8) score. - (’film.film.language’, ’English’)is not related to any information in the sentence,(0) score. sentence: {} information in the sentence: {} candidate entity: {} assess:
answering please answer the question based on the provided triplets. The format of the triplet is (entity1, relation, entity2). If there is not enough information provided in the triplets, answer the question with your own knowledge. question: Which place is the madam satan located? triplets: [(’madam satan’, ’film.film.country’, ’the USA’), (’madam satan’, ’film.film.language’, ’English’)] answer: ’madam satan’ is located in ’the USA’. question: {} triplets: {} answer:
Figure 3: LLM prompts for Chinese datasets.
A.2 Pseudocode
The pseudocode of Adaptive and Stateful Knowledge Exploration can be seen in Algorithm 1.
Input: Initial entity $e^{0}$, all clues $clue_{all}$, knowledge graph $\mathcal{G}$
Output: Knowledge path $P$, explored clues $C$
Initialize $P\leftarrow\emptyset$, $C\leftarrow\emptyset$, $n\leftarrow 1$
$C^{\prime}\leftarrow clue_{all}\setminus C$
  // Candidate clues
while not terminated do
       // Relation-Based Mapping (Step $n$)
       $R\leftarrow\text{GetRelations}(\mathcal{G},e^{n-1})$
[…]
             Terminate
       end if
      // Entity-Based Mapping (Step $n$)
[…]
             $E_{j}\leftarrow\text{GetEntities}(\mathcal{G},e^{n-1},r_{j})$
       end foreach
[…]
             Terminate
              // Case 3: No entity mapping
      else
             // Handle successful mappings
             if $\exists(\text{clue}_{1},r_{i}),(\text{clue}_{1},e_{j})$ then
                    // Case 1: Single-clue dual mapping
                   $C\leftarrow C\cup\{\text{clue}_{1}\}$
                   $P\leftarrow P\cup\{(e^{n-1},r_{i},e_{j})\}$
                   $e^{n}\leftarrow e_{j}$
             end if
            else if $\exists(\text{clue}_{1},r_{i}),(\text{clue}_{2},e_{j})$ then
                    // Case 2: Dual-clue mapping
                   $C\leftarrow C\cup\{\text{clue}_{1},\text{clue}_{2}\}$
                   $P\leftarrow P\cup\{(e^{n-1},r_{i},e_{j})\}$
                   $e^{n}\leftarrow e_{j}$
             end if
            $n\leftarrow n+1$
             $C^{\prime}\leftarrow clue_{all}\setminus C$
       end if
end while
return $P,C$
Algorithm 1 Pseudocode of Adaptive and Stateful Knowledge Exploration
[…]